OpenStudy (anonymous):

Solve this system of equations ; 3/4x+ 1/2y = -3/4 2/3x + 2/3y = 2/3

OpenStudy (anonymous):

Get rid of the fractions. Muliply both sides of the first equation by 4 and do the same to the second equation using 3.

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

\[4\left(\frac{3 x}{4}+\frac{y}{2}\right)=-\frac{3}{4}*4\text{ }\to \text{ }3 x+2 y=-3\]\[3\left(\frac{2 x}{3}+\frac{2 y}{3}\right)=\frac{2}{3}*3\text{ }\to \text{ }2x+2y=2\]

OpenStudy (anonymous):

so , is that the answer ?

OpenStudy (anonymous):

Once the fractions are gone, then you have two equations in two unknowns, x and y. {x = -5, y = 6} Can you solve the equations?

OpenStudy (anonymous):

Yes , Thanks.

OpenStudy (anonymous):

You can verify the values of x and y by plugging them into the original equations with the fractions.

OpenStudy (anonymous):

Thank you for the medal.